# These $\equalcontThese authors contributed equally to this work.$

### v2

@@
[…]
These $\equalcontThese authors contributed equally to this work.$
[…]
@@
[…]
keywords:
@@
Sequential recommendations (SR) with transformer-based architectures are widely adopted in real-world applications, where SR models require frequent retraining to adapt to ever-changing user preferences. However, training transformer-based SR models often encounters a high computational cost associated with scoring extensive item catalogs, often exceeding thousands of items. This occurs mainly due to the use of cross-entropy loss, where peak memory scales proportionally to catalog size, batch size, and sequence length. Recognizing this, practitioners in the field of recommendation systems typically address memory consumption by integrating the cross-entropy (CE) loss with negative sampling, thereby reducing the explicit memory demands of the final layer. However, a small number of negative samples would degrade model performance, and as we demonstrate in our work, increasing the number of negative samples and the batch size further improves the model’s performance, but rapidly starts to exceed even industrial GPUs’ size ( 40Gb). While it is possible to address this problem with multi-GPU settings, it increases costs, and the scaling of training time becomes a concern due to inter-device communication.
In this work, we introduce the $\mathrm{CCE}^{-}$ method, which offers a GPU-efficient implementation of the CE loss with negative sampling. As a result, our method accelerates training by up to two times while reducing memory consumption by more than 10 times. Leveraging the memory savings afforded by using $\mathrm{CCE}^{-}$ for model training, it becomes feasible to enhance its accuracy on datasets with a large item catalog compared to those trained with original PyTorch-implemented loss functions. Finally, we perform an analysis of key memory-related hyperparameters and highlight the necessity of a delicate balance among these factors. For example, we demonstrate that scaling both the number of negative samples and batch size leads to better results rather than maximizing only one of them. To facilitate further adoption of $\mathrm{CCE}^{-}$, we release a Triton kernel that efficiently implements the proposed method. [^1] Code, reproducibility materials, and all scripts for generating figures are available at https://github.com/On-Point-RND/MemoryEfficientSRS.
1 Introduction
User actions can be modeled sequentially, conditioning on past actions in diverse scenarios, ranging from a mechanical series of actions in a banking app [4] to a mood- and neighborhood-influenced selection of music tracks [18]. Consequently, Sequential Recommender Systems (SRS), which model user interactions as sequences, have become essential components of modern recommendation pipelines. SRS effectively incorporates ideas from natural language processing (NLP) through attention-based architectures such as SASRec [15] and BERT4Rec [29], as well as the cross-entropy loss function [17, 29, 21].
@@ 1 Introduction
Peak memory is one of the main computational constraints when cross-entropy loss (CE) is applied to large catalogs. The memory footprint at the final layer for logits can be estimated as $bs\cdot sl\cdot|V|$ [^1] To illustrate, consider a batch of 256 sequences, each of length 100, with a catalog containing $10^{6}$ items. The resulting logit tensor would require the storage of $256\cdot 100\cdot 10^{6}$ elements, equating to approximately 100 GB of memory., where $bs$ is the batch size, $sl$ sequence length and $|V|$ is the catalog or vocabulary size, typically the most significant contributor. This final layer uses up to 90% of the combined memory footprint, gradients, and activations, when training modern Large Language Models (LLMs) [36]. However, in SRS, $|V|$ can reach millions of items [41], surpassing even LLMs, which have vocabulary sizes orders of magnitude smaller: 32K, 152K, and 256K for Llama2-7B [32], QWEN [3], and Gemma-7B [30], respectively.
Figure 1: Here, we compare several methods CE, $\mathrm{CE}^{-}$, CCE, $\mathrm{CCE}^{-}$, and SCE in terms of (1) NDCG@10, (2) Training time per epoch in seconds, and (3) Memory consumption in Gb for the SASRec model across six datasets. We demonstrate the best performance achieved through optimized hyperparameters. We can see that CE, $\mathrm{CE}^{-}$ require much more time and memory usage.
Table 1: NDCG@10 metric values for the SASRec model trained using different loss functions. The best metric value is in bold, and the second-best is underlined.
@@
This memory footprint has been a key driver for early SRS adopting the binary cross-entropy loss (BCE) [15] and the cross-entropy loss with negative sampling ($\mathrm{CE}^{-}$[29, 16]. $\mathrm{CE}^{-}$ directly addresses memory constraints caused by large item catalogs [16, 21]. Instead of scoring all items during training, a subset of negative samples ($ns$) is used, where $ns\ll|V|$, to approximate the full CE loss. This reduces computational complexity and memory requirements but introduces a hyperparameter trade-off between batch size, sequence length, and number of negative samples (ns) in memory-constrained scenarios. Prior work indicates that maximizing $ns$ in $\mathrm{CE}^{-}$ or using CE can improve performance of the models [8, 16, 9]. However, empirical results demonstrate that $\mathrm{CE}^{-}$ can sometimes outperform full CE for some datasets [21, 16] and the performance of models trained with $\mathrm{CE}^{-}$ shows a non-linear dependence on nsns[24]. This creates an open question of whether practitioners should maximize $ns$ given a fixed memory budget or prioritize other memory-related parameters. We will refer to this question as the first question Q1.
@@
Although sampling strategies such as $\mathrm{CE}^{-}$ reduce memory consumption, they also improve training speed by requiring fewer operations, partially addressing the problem of frequent retraining SRS. However, there are other approaches for faster training: quantization [37], activation sparsification [14], and finally gradient sparsification [1]. The latter one is usually overlooked, even though gradients usually require twice as many computations as the forward path. In our work, we demonstrate that gradient sparsification is especially suitable for SRS, provide a formal analysis in Section 3.2, and study how sparse gradients accelerate model training and their impact on accuracy degradation. We will refer to it as our second research question Q2.
@@
Computational efficiency critically depends on the co-design of architecture, hardware, and low-level implementations. While hardware is often fixed, efficiency can be optimized through architectures better aligned with underlying hardware [27] or specialized low-level optimizations. Though highly specific, such optimizations yield significant computational benefits. Examples include FlashAttention [7] for efficient attention computation and Liger kernels [13] for optimized CE computation. Among them, Cut-Cross Entropy (CCE) was recently developed for LLMs; it helps to alleviate memory bottlenecks during training when CEloss is employed. However, until now, the CCE has remained unexplored for SRS models. Furthermore, our analysis reveals the importance of negative sampling, which is not supported by the original CCE. Development and analysis of CCE with negative sampling is our third research question - Q3, more precisely: Can we design a CCE alternative that is faster and more memory efficient for SRS?
@@
  1. 1.
    Comprehensive Memory-Accuracy Trade-off Analysis: To address Q1, we conducted a systematic investigation into the relationships among accuracy of SASRec, batch size ($bs$), sequence length ($sl$), the number of negative sample size ($ns$) across six datasets with varying item catalog sizes ($|V|$). Our findings indicate that simply maximizing $ns$ in $\mathrm{CE}^{-}$ or employing CE to encompass the entire item catalog $|V|$ leads to suboptimal performance in memory-constrained environments. Instead, the choice between $\mathrm{CE}^{-}$ and CE requires a data set-specific calibration. We provide analysis for configuring these parameters to balance accuracy and memory efficiency.
    Furthermore, we make publicly available a dataset comprising 997 data points, each representing SASRec trained across the considered datasets. For every configuration, we log $bs$, slsl, nsns, memory consumption, and various metrics along the training steps. This dataset enables reproducible analysis of memory-accuracy trade-offs.
  2. 2.
@@
    Memory-Efficient Metric Improvement: By reallocating the memory savings achieved through the use of CCE and $\mathrm{CCE}^{-}$ to increase batch size $bs$, sequence length $sl$, and the number of negative samples $ns$, we attain up to a 30% improvement in accuracy while maintaining the same computational budget.
2 Related work
@@
Cross-Entropy and sampled CE for SRS. Sequential recommendation has seen significant advancements in recent years. One pivotal observation from  [16] demonstrated that SASRec’s performance improves when trained with cross-entropy loss instead of binary cross-entropy. Unfortunately, computing cross-entropy over all items is computationally prohibitive. To address this, researchers adopted sampled cross-entropy loss, which uses a subset of negative items. This approach was theoretically justified in  [8], showing that increasing negative samples enhances performance. However, negative sampling has non-linear performance scaling. Despite its benefits, negative sampling amplifies popularity bias, as popular items are more likely to be sampled  [24]. Consequently, model performance becomes highly dependent on dataset distribution and sampling strategy. Furthermore,  [24] highlights that performance does not scale linearly with the number of negative samples, which complicates the selection of hyperparameters. To address this problem, we perform a careful and extensive study of memory-dependent hyperparameter analysis for optimal accuracy-memory tradeoff (Section 5.1). Prior work on scaling for SRS models  [9, 39, 41, 38] emphasizes the benefits of scaling sequence length and model depth. However, our experiments reveal that expanding a single factor [^3] Factors include batch size, sequence length, and number of negative samples. yields limited gains, underscoring the need for multi-dimensional scaling. Unlike  [21], our work focuses only on feature-scaling [^4] Feature scaling refers to adjusting batch size, sequence length, etc. effects while keeping model architecture fixed.
@@
Computationally efficient approaches. Advanced sampling techniques for balancing accuracy and memory were proposed in  [21]. Liger-Kernel  [13], implemented in the Triton language  [31], addresses memory consumption issues through the use of FusedLinearCrossEntropy (FLCE). While theoretical convergence guarantees for sparse gradient methods were established in  [1], it’s application for efficient training remains limited  [5, 36]. Building on FLCE memory-efficient cross-entropy with saturated gradient filtering beyond numerical precision was proposed (CCE) for LLMs  [36]. Since saturated gradients do not fit numerical precision and turn to zero, their removal does not affect model convergence. In this work, we analyze whether the removal of low-magnitude non-zero gradients would help to accelerate model training with minimal performance degradation.
@@
Possible further directions and trends. Even when isolating experiments to feature scaling, there remains significant flexibility in training strategies. For example, work  [19] demonstrated that training GPT-like models with short sequences initially and gradually increasing them enables stable, efficient training with larger batch sizes and learning rates. Similarly, authors of  [8] found that starting with fewer negative samples and increasing their count later boosts SRS model performance. Moreover, based on our results, an adaptive threshold to prune gradients can be employed. While these adaptive strategies are promising, we focus on studying simpler non-adaptive approaches to establish a stronger foundation for future research.
3 Methods
@@
Figure 2: Illustration of CE, CCE and $\mathrm{CCE}^{-}$ implementation. CE: Materializes full logits in GPU memory, matmul and softmax are computed separately $\rightarrow$ high GPU memory usage, full $C\times E$ matmul, extra data transfers between HBM and SRAM lead to higher time delay. CCE : No logits materialization, fused matmul + softmax $\rightarrow$ store only positive logits and LSE vector, reduced HBM $\leftarrow\rightarrow$ SRAM transfers, but still full CxE matmul. $\mathrm{CCE}^{-}$: No logits materialization, fused matmul + softmax on sampled items $\rightarrow$ store only positive logits and LSE vector, reduced HBM $\leftarrow\rightarrow$ SRAM transfers, matmul only for sampled columns of C and corresponding E.
3.1 Cut-Cross-Entropy for RecSys
@@
Our proposed method, $\mathrm{CCE}^{-}$, builds upon CCE, which itself extends the Liger-Kernel (LG) [13]. Below, we provide a brief review of both $\mathrm{CCE}^{-}$ and LG, highlighting their key differences. We then focus on how $\mathrm{CCE}^{-}$ introduces specific improvements tailored for sequential recommendation systems.
@@
During training, we deal with two types of memory: GPU memory (High Bandwidth Memory, HBM) and on-chip shared memory (Static Random-Access Memory, SRAM). HBM is large but relatively slow and is used to store data for computations. GPU kernel loads input data from HBM into SRAM to perform computations. SRAM is significantly faster than HBM but has much less capacity. Therefore, only small chunks of input data can be loaded into SRAM at a time. Once computations are complete, the output is saved back to HBM. While there are more details, see [7],  [36], this structure shapes the main goal: reduce the amount of data loaded into GPU memory. It can be achieved by minimizing the overall amount of memory used by a model or optimizing transfers between HBM and SRAM.
@@
In this study, we incorporated CCE into sequential recommendation transformer models to evaluate its effectiveness for recommendation tasks. For this purpose, the three-dimensional output tensor of embeddings produced by the final transformer block is flattened into a matrix $\mathbf{E}\in\mathbb{R}^{N\times D}$, where $D$ represents the hidden size and $N$ corresponds to the total number of items, calculated as the product of the batch size $bs$ and the sequence length $sl$. The CE loss for the set of items of size |Vvert|V\ vert is written as follows:
$\mathcal{L}$ $=-\dfrac{1}{N}\sum_{i=1}^{N}\mathbf{l}_{i},$ (1)
$\mathbf{l}$ $=(\mathbf{C}^{\mathrm{T}}\mathbf{E})_{\mathbf{x}}-\mathrm{log}\sum_{j=1}^{V}\exp{(\mathbf{C}^{\mathrm{T}}\mathbf{E})},$ (2)
@@
Figure 3: An illustration of computing gradients for the weights of the final layer.
3.2 Sparse Gradients in Training with CE loss
In the recommender systems, the objective is to predict the next item from a predefined set $V$ of |V||V|{} items. Similar to the next-token prediction, there is exactly one correct class and |V|1|V|{}-1 incorrect classes. Thus, employing loss functions such as the CE loss provides a suitable estimate of model quality  [8].

### v3

# Faster and Memory-Efficient Training of Sequential Recommendation Models for Large Catalogs

@@
Maxim Zhelnin zhelninmax@gmail.com MWS AIMoscowRussia , Dmitry Redko dmitryredko444@gmail.com Applied AIMoscowRussia , Daniil Volkov danvol121@gmail.com Applied AIMoscowRussia , Anna Volodkevich volodkanna@yandex.ru Sber AI Lab, Applied AIMoscowRussia , Petr Sokerin sokerinpo@gmail.com Applied AIMoscowRussia , Valeriy Shevchenko escape756@gmail.com IVIMoscowRussia , Egor Shvetsov e.shvetsov@applied-ai.ru Applied AIMoscowRussia , Alexey Vasilev alexxl.vasilev@yandex.ru Sber AI LabMoscowRussia , Darya Denisova duny.explorer@gmail.com Sber AI LabMoscowRussia , Ruslan Izmailov lord.rik@yandex.ru SberMoscowRussia and Alexey Zaytsev likzet@gmail.com Applied AIMoscowRussia
@@
$c$s: Information systems Recommender systems
Sequential recommendations (SR) with transformer-based architectures are widely adopted, and SR models require frequent retraining to adapt to changing user preferences. However, training transformer-based SR models incurs high computational cost from scoring large item catalogs (often thousands of items). This is mainly due to cross-entropy loss, where peak memory scales with catalog size, batch size, and sequence length. To reduce memory, practitioners combine cross-entropy (CE) with negative sampling, lowering explicit memory needs of the final layer. Yet few negative samples degrade performance; increasing negatives and batch size improves results but quickly exceeds industrial GPU memory ( 40Gb).
In this work, we introduce $\mathrm{CCE}^{-}$, a GPU-efficient implementation of CE with negative sampling. $\mathrm{CCE}^{-}$ accelerates training by up to 2× and reduces memory by more than 10×. The saved memory enables higher accuracy on datasets with large catalogs versus models trained with standard PyTorch losses. We also analyze key memory-related hyperparameters and show the need to balance them: scaling both the number of negative samples and batch size yields better results than maximizing only one. To support adoption, we release a Triton kernel implementing the method efficiently. [^1] Code, reproducibility materials, and all scripts for generating figures are available at https://github.com/On-Point-RND/MemoryEfficientSRS
1. Introduction
User actions can be modeled sequentially, conditioning on past actions in diverse scenarios, ranging from a mechanical series of actions in a banking app (boka2024survey) to a mood- and neighborhood-influenced selection of music tracks (krause2014music). Consequently, Sequential Recommender Systems (SRS), which model user interactions as sequences, have become essential components of modern recommendation pipelines. SRS effectively incorporates ideas from natural language processing (NLP) through attention-based architectures such as SASRec (kang2018self) and BERT4Rec (sun2019bert4rec), as well as the cross-entropy loss function (klenitskiy2024does; sun2019bert4rec; mezentsev2024scalable).
@@ 1. Introduction
Peak memory is one of the main computational constraints when cross-entropy loss (CE) is applied to large catalogs. The memory footprint at the final layer for logits can be estimated as $bs\cdot sl\cdot|V|$ [^1] To illustrate, consider a batch of 256 sequences, each of length 100, with a catalog containing $10^{6}$ items. The resulting logit tensor would require the storage of $256\cdot 100\cdot 10^{6}$ elements, equating to approximately 100 GB of memory., where bsbs is the batch size, slsl sequence length and $|V|$ is the catalog or vocabulary size, typically the most significant contributor. This final layer uses up to 90% of the combined memory footprint, gradients, and activations, when training modern Large Language Models (LLMs) (wijmans2024cut). However, in SRS, $|V|$ can reach millions of items (zivic2024scaling), surpassing even LLMs, which have vocabulary sizes orders of magnitude smaller: 32K, 152K, and 256K for Llama2-7B (touvron2023llama), QWEN (bai2023qwen), and Gemma-7B (team2403gemma), respectively.
Figure 1. Here, we compare several methods CE, $\mathrm{CE}^{-}$, CCE, $\mathrm{CCE}^{-}$, and SCE in terms of (1) NDCG@10, (2) Training time per epoch in seconds, and (3) Memory consumption in Gb for the SASRec model across six datasets. We demonstrate the best performance achieved through optimized hyperparameters. We can see that CE, $\mathrm{CE}^{-}$ require much more time and memory usage.
Table 1. NDCG@10 metric values for the SASRec model trained using different loss functions. The best metric value is in bold, and the second-best is underlined.
@@
This memory footprint has been a key driver for early SRS adopting the binary cross-entropy loss (BCE) (kang2018self) and the cross-entropy loss with negative sampling ($\mathrm{CE}^{-}$(sun2019bert4rec; klenitskiy2023turning). $\mathrm{CE}^{-}$ directly addresses memory constraints caused by large item catalogs (klenitskiy2023turning; mezentsev2024scalable). Instead of scoring all items during training, a subset of negative samples (nsns) is used, where $ns\ll|V|$, to approximate the full CE loss. This reduces computational complexity and memory requirements but introduces a hyperparameter trade-off between batch size, sequence length, and number of negative samples (ns) in memory-constrained scenarios. Prior work indicates that maximizing nsns in $\mathrm{CE}^{-}$ or using CE can improve performance of the models (di2024theoretical; klenitskiy2023turning; guo2024scaling). However, empirical results demonstrate that $\mathrm{CE}^{-}$ can sometimes outperform full CE for some datasets (mezentsev2024scalable; klenitskiy2023turning) and the performance of models trained with $\mathrm{CE}^{-}$ shows a non-linear dependence on nsns(prakash2024evaluating). This creates an open question of whether practitioners should maximize nsns given a fixed memory budget or prioritize other memory-related parameters. We will refer to this question as the first question Q1.
@@
Although sampling strategies such as $\mathrm{CE}^{-}$ reduce memory consumption, they also improve training speed by requiring fewer operations, partially addressing the problem of frequent retraining SRS. However, there are other approaches for faster training: quantization (wortsman2023stable), activation sparsification (hu2024accelerating), and finally gradient sparsification (alistarh2018convergence). The latter one is usually overlooked, even though gradients usually require twice as many computations as the forward path. In our work, we demonstrate that gradient sparsification is especially suitable for SRS, provide a formal analysis in Section 3.2, and study how sparse gradients accelerate model training and their impact on accuracy degradation. We will refer to it as our second research question Q2.
@@
Computational efficiency critically depends on the co-design of architecture, hardware, and low-level implementations. While hardware is often fixed, efficiency can be optimized through architectures better aligned with underlying hardware (shvetsov2024quantnas) or specialized low-level optimizations. Though highly specific, such optimizations yield significant computational benefits. Examples include FlashAttention (dao2022flashattention) for efficient attention computation and Liger kernels (hsu2024liger) for optimized CE computation. Among them, Cut-Cross Entropy (CCE) was recently developed for LLMs; it helps to alleviate memory bottlenecks during training when CEloss is employed. However, until now, the CCE has remained unexplored for SRS models. Furthermore, our analysis reveals the importance of negative sampling, which is not supported by the original CCE. Development and analysis of CCE with negative sampling is our third research question - Q3, more precisely: Can we design a CCE alternative that is faster and more memory efficient for SRS?
@@
  1. (1)
    Comprehensive Memory-Accuracy Trade-off Analysis: To address Q1, we conducted a systematic investigation into the relationships among accuracy of SASRec, batch size (bsbs), sequence length (slsl), the number of negative sample size (nsns) across six datasets with varying item catalog sizes ($|V|$). Our findings indicate that simply maximizing nsns in $\mathrm{CE}^{-}$ or employing CE to encompass the entire item catalog $|V|$ leads to suboptimal performance in memory-constrained environments. Instead, the choice between $\mathrm{CE}^{-}$ and CE requires a data set-specific calibration. We provide analysis for configuring these parameters to balance accuracy and memory efficiency.
    Furthermore, we make publicly available a dataset comprising 997 data points, each representing SASRec trained across the considered datasets. For every configuration, we log bsbs, slsl, nsns, memory consumption, and various metrics along the training steps. This dataset enables reproducible analysis of memory-accuracy trade-offs.
  2. (2)
@@
    Memory-Efficient Metric Improvement: By reallocating the memory savings achieved through the use of CCE and $\mathrm{CCE}^{-}$ to increase batch size bsbs, sequence length slsl, and the number of negative samples nsns, we attain up to a 30% improvement in accuracy while maintaining the same computational budget.
2. Related work
@@
Cross-Entropy and sampled CE for SRS. Sequential recommendation has seen significant advancements in recent years. One pivotal observation from  (klenitskiy2023turning) demonstrated that SASRec’s performance improves when trained with cross-entropy loss instead of binary cross-entropy. Unfortunately, computing cross-entropy over all items is computationally prohibitive. To address this, researchers adopted sampled cross-entropy loss, which uses a subset of negative items. This approach was theoretically justified in  (di2024theoretical), showing that increasing negative samples enhances performance. However, negative sampling has non-linear performance scaling. Despite its benefits, negative sampling amplifies popularity bias, as popular items are more likely to be sampled  (prakash2024evaluating). Consequently, model performance becomes highly dependent on dataset distribution and sampling strategy. Furthermore,  (prakash2024evaluating) highlights that performance does not scale linearly with the number of negative samples, which complicates the selection of hyperparameters. To address this problem, we perform a careful and extensive study of memory-dependent hyperparameter analysis for optimal accuracy-memory tradeoff (Section 5.1). Prior work on scaling for SRS models  (guo2024scaling; zhang2024scaling; zivic2024scaling; xu2025efficient) emphasizes the benefits of scaling sequence length and model depth. However, our experiments reveal that expanding a single factor [^3] Factors include batch size, sequence length, and number of negative samples. yields limited gains, underscoring the need for multi-dimensional scaling. Unlike  (mezentsev2024scalable), our work focuses only on feature-scaling [^4] Feature scaling refers to adjusting batch size, sequence length, etc. effects while keeping model architecture fixed.
@@
Computationally efficient approaches. Advanced sampling techniques for balancing accuracy and memory were proposed in  (mezentsev2024scalable). Liger-Kernel  (hsu2024liger), implemented in the Triton language  (tillet2019triton), addresses memory consumption issues through the use of FusedLinearCrossEntropy (FLCE). While theoretical convergence guarantees for sparse gradient methods were established in  (alistarh2018convergence), it’s application for efficient training remains limited  (chekalina2024sparsegrad; wijmans2024cut). Building on FLCE memory-efficient cross-entropy with saturated gradient filtering beyond numerical precision was proposed (CCE) for LLMs  (wijmans2024cut). Since saturated gradients do not fit numerical precision and turn to zero, their removal does not affect model convergence. In this work, we analyze whether the removal of low-magnitude non-zero gradients would help to accelerate model training with minimal performance degradation.
@@
Possible further directions and trends. Even when isolating experiments to feature scaling, there remains significant flexibility in training strategies. For example, work  (li2022stability) demonstrated that training GPT-like models with short sequences initially and gradually increasing them enables stable, efficient training with larger batch sizes and learning rates. Similarly, authors of  (di2024theoretical) found that starting with fewer negative samples and increasing their count later boosts SRS model performance. Moreover, based on our results, an adaptive threshold to prune gradients can be employed. While these adaptive strategies are promising, we focus on studying simpler non-adaptive approaches to establish a stronger foundation for future research.
3. Methods
@@
Figure 2. Illustration of CE, CCE and $\mathrm{CCE}^{-}$ implementation. CE: Materializes full logits in GPU memory, matmul and softmax are computed separately $\rightarrow$ high GPU memory usage, full $C\times E$ matmul, extra data transfers between HBM and SRAM lead to higher time delay. CCE : No logits materialization, fused matmul + softmax $\rightarrow$ store only positive logits and LSE vector, reduced HBM $\leftarrow\rightarrow$ SRAM transfers, but still full CxE matmul. $\mathrm{CCE}^{-}$: No logits materialization, fused matmul + softmax on sampled items $\rightarrow$ store only positive logits and LSE vector, reduced HBM $\leftarrow\rightarrow$ SRAM transfers, matmul only for sampled columns of C and corresponding E.
3.1. Cut-Cross-Entropy for RecSys
@@
Our proposed method, $\mathrm{CCE}^{-}$, builds upon CCE, which itself extends the Liger-Kernel (LG) (hsu2024liger). Below, we provide a brief review of both $\mathrm{CCE}^{-}$ and LG, highlighting their key differences. We then focus on how $\mathrm{CCE}^{-}$ introduces specific improvements tailored for sequential recommendation systems.
@@
During training, we deal with two types of memory: GPU memory (High Bandwidth Memory, HBM) and on-chip shared memory (Static Random-Access Memory, SRAM). HBM is large but relatively slow and is used to store data for computations. GPU kernel loads input data from HBM into SRAM to perform computations. SRAM is significantly faster than HBM but has much less capacity. Therefore, only small chunks of input data can be loaded into SRAM at a time. Once computations are complete, the output is saved back to HBM. While there are more details, see (dao2022flashattention),  (wijmans2024cut), this structure shapes the main goal: reduce the amount of data loaded into GPU memory. It can be achieved by minimizing the overall amount of memory used by a model or optimizing transfers between HBM and SRAM.
@@
In this study, we incorporated CCE into sequential recommendation transformer models to evaluate its effectiveness for recommendation tasks. For this purpose, the three-dimensional output tensor of embeddings produced by the final transformer block is flattened into a matrix $\mathbf{E}\in\mathbb{R}^{N\times D}$, where $D$ represents the hidden size and NN corresponds to the total number of items, calculated as the product of the batch size bsbs and the sequence length slsl. The CE loss for the set of items of size |Vvert|V\ vert is written as follows:
(1) $\mathcal{L}$ $=-\dfrac{1}{N}\sum_{i=1}^{N}\mathbf{l}_{i},$
(2) $\mathbf{l}$ $=(\mathbf{C}^{\mathrm{T}}\mathbf{E})_{\mathbf{x}}-\mathrm{log}\sum_{j=1}^{V}\exp{(\mathbf{C}^{\mathrm{T}}\mathbf{E})},$
@@
Figure 3. An illustration of computing gradients for the weights of the final layer.
3.2. Sparse Gradients in Training with CE loss
In the recommender systems, the objective is to predict the next item from a predefined set $V$ of |V||V|{} items. Similar to the next-token prediction, there is exactly one correct class and |V|1|V|{}-1 incorrect classes. Thus, employing loss functions such as the CE loss provides a suitable estimate of model quality  (di2024theoretical).